# Vision-Language Models for Medical Report Generation and Visual Question Answering: A Review

### v1

@@ 5.2 Evaluation Metrics for Visual Question Answering
  • Human evaluation is valuable for assessing a model’s performance and applies not only to tasks such as VQA but also to RG. Human evaluation can be performed for VQA in various ways. For instance, in [Moo+23], the human evaluation process of MedFlamingo model employs an application featuring a user-friendly interface. Within this interface, medical experts are empowered to evaluate each VQA problem individually, assigning scores ranging from 0 to 10. The final scores of the few-shot performance are $5.61$ on VQA-RAD, $1.81$ on PathVQA, and $4.33$ on the specifically curated Visual USMLE dataset. In contrast, the scores for zero-shot performance are lower, with $3.82$ on RAD-VQA, $1.72$ on PathVQA, and $4.18$ on Visual USMLE.
@@
[Moo+22] Open-I, VQA-RAD
PubMedCLIP dual No ViT-B/32 or RN50 + ROCO, SLAKE, GH
[EMD23] or RN50$\times$4 VQA-RAD
+ Transformer + BAN
RepsNet dual No ResNeXt-101 + BERT + + VQA-RAD, IU-Xray Site
[TBF22] + BAN + GPT-2
@@
UniXGen single No VQGAN + MIMIC-CXR GH
[Lee+23] + Transformer
@@
[Yua+23] + multimodal encoder SLAKE, VQA-RAD,
w/ retrieval-attention VQA-Med 2019,
module VQA-Med 2021
X-REM dual No ALBEF + MIMIC-CXR, GH
[Jeo+23] (ViT-B/16 + BERT MedNLI, RadNLI
+ multimodal encoder)
@@
[Shu+23] +LLaMa-7B MedQA, PubMedQA
@@
CXR-RePaiR-Gen + prompt manager
[Ran+23a] + text-davinci-003
@@
LLaVa-Med single No ViT-L/14 + PMC-15 + GPT-4, GH
[Li+23] + projection layer VQA-RAD, SLAKE,
+ LLaMa-7B PathVQA
XrayGPT single No MedCLIP + linear + + MIMIC-CXR, GH
[Tha+23] transformation layer Open-I
+ Vicuna-7B
@@
dual Yes ViT-B/12 + BERT + ROCO, MedICaT, GH
MUMC + multimodal encoder ImageCLEF Caption,
[Li+23a] + answer decoder VQA-RAD, SLAKE
PathVQA
Med-Flamingo single No ViT-L/14 + MTB, PMC-OA, GH
[Moo+23] + perceiver resampler VQA-RAD, PathVQA,
+ LLaMa-7B Visual USMLE
RaDialog single No BioViL-T + BERT + + MIMIC-CXR, GH
[Pel+23] + prompt manager Instruct
+ Vicuna-7B
Table 2: A list of medical VLMs designed for VQA and/or RG.
6 Medical Vision-Language Models
@@
This section presents an overview of existing medical VLMs tailored for visual question answering and/or report generation. The information is organized chronologically based on their first appearance. Our focus is mainly on models with available code, primarily those introduced more recently. A list of these VLMs is presented in Table 2.
@@
  • MedViLL (Medical Vision Language Learner) [Moo+22] is a VLM specifically developed for handling medical images and generating associated reports. It employs ResNet-50 [He+16], trained on ImageNet [Den+09], for extracting visual feature vector $\vec{v}$. The model also leverages a base BERT [Dev+19] embedding layers to extract textual feature vectors $\vec{t}$ from clinical reports, which are initially segmented into a sequence of tokens using a WordPiece [Wu+16] tokenizer. Both textual and visual features incorporate positional information to capture the spatial relationships and sequential order of elements in the input data. To generate a cross-modal representation, vectors $\vec{v}$ and $\vec{t}$, along with special tokens [CLS] and [SEP], are concatenated as follows: $(\overrightarrow{CLS},\vec{v},\overrightarrow{SEP_{V}},\vec{t},\overrightarrow{SEP_{L}})$. The cross-modal representations are then fed into the BERT model. The MedViLL is pre-training on two tasks: MLM and ITM. The MLM task employs a bidirectional auto-regressive (BAR) self-attention mask, promoting the integration of image and language features. For MLM, a negative log-likelihood loss function is used. The ITM task encourages learning visual and textual features by predicting matching pairs and employs a loss function based on predictions for matching and non-matching pairs. The model is pre-trained on $89,395$ image-report pairs from the MIMIC-CXR [Joh+19a] dataset and then fine-tuned for downstream tasks on $3,547$ pairs from the Open-I [Dem+15] dataset, an additional dataset comprising radiographic image-report pairs. Only AP view X-rays are included from both datasets for the analysis. VQA is performed on the VQA-RAD [Lau+18] dataset (see Table 3), including open and close-ended questions, where the output representation of [CLS] is used to predict a one-hot encoded answer. For radiology RG fine-tuning, the model uses sequence-to-sequence (S2S) mask instead of BAR and generates reports by sequentially recovering MASK tokens. RG is evaluated on MIMIC-CXR [Joh+19a] and Open-I [Dem+15]. MedViLL achieves a BLEU-4 score of $0.066$, a perplexity value of $4.185$, and using a CheXpert labeler [Irv+19] an accuracy of $84.1\%$, a precision value of $0.698$, a recall value of $0.559$, and an F1 score of $0.621$ on MIMIC-CXR. Additionally, it achieves a BLEU-4 score of $0.049$, a perplexity value of $5.637$, an accuracy of $73.4\%$, a precision value of $0.512$, a recall value of $0.594$, and an F1 score of $0.550$ on Open-I.
@@
  • PubMedCLIP [EMD23] is a CLIP-based [Rad+21] model pre-trained on ROCO [Pel+18] dataset, consisting of over 80K image-caption pairs sourced from PMC articles. The model utilizes CLIP text encoder, which is based on a Transformer [Vas+17] architecture, and three distinct CLIP visual encoders: ViT-B/32 [Dos+21], ResNet-50, and ResNet-50$\times$4 [He+16]. Following the contrastive learning approach in CLIP, the model generates joint representations by computing cosine similarity between textual and visual features. The pre-training objective involves the computation of cross-entropy loss values for both vision and language. These losses are then averaged to derive an overall loss value. Following pre-training, the model is repurposed as a pre-trained visual encoder for VQA. The visual feature in VQA is the concatenation of the model’s output with the convolutional denoising autoencoder (CDAE) [Mas+11] output, an image denoising module. The question is encoded using a GloVe [PSM14] word embedding followed by an LSTM [HS97]. The image and question features are combined using bilinear attention networks (BAN) [KJZ18], and the resulting representations are passed through an answer classifier, which is a two-layer feedforward NN. The VQA loss is determined by combining the classification and image reconstruction losses. During the VQA fine-tuning, the SLAKE (English) [Liu+21] and VQA-RAD [Lau+18] datasets, comprising both open- and close-ended questions, are employed. The model’s effectiveness is evaluated in the context of two existing MedVQA (Medical Visual Question Answering) methods: MEVF (Mixture of Enhanced Visual Features) [Zha+20] and QCR question-conditioned reasoning (QCR) [Liu+23]. The assessment involved replacing the visual encoder component in MEVF and QCR with PubMedCLIP and subsequently evaluating the model’s performance. PubMedCLIP in the QCR framework achieves better accuracies on VQA-RAD and SLAKE datasets than in the MEVF framework. The highest accuracies of PubMedCLIP in the QCR framework on both datasets are shown in Table 3.
@@
  • RepsNet [TBF22] model is designed in a VQA setting and can generate automated medical reports and interpret medical images. It employs a modified version of the pre-trained ResNeXt-101 [Xie+16] as its image encoder and utilizes pre-trained BERT [Dev+19] as the text encoder, with text tokenization done through WordPiece [Wu+16]. Fusion of image and question features is achieved using BAN [KJZ18]. To align images with textual descriptions, the model employs bidirectional contrastive learning [Che+20]. For VQA tasks, the model undergoes fine-tuning and evaluation on VQA-RAD [Lau+18] (see Table 3), while for RG, fine-tuning and evaluation take place on IU-Xray [Dem+15] dataset. The model categorizes answers through classification for close-ended questions and generates answers using the modified version of GPT-2 language decoder based on image features and prior context. The BLEU-2 and BLEU-4 scores of RepsNet on the IU-Xray dataset are $0.44$ and $0.27$, respectively.
@@
  • BiomedCLIP [Zha+23a] is pre-trained on the specifically curated PMC-15 dataset that consists of $15$ M figure-caption pairs derived from the PMC articles, however, it is not publicly available. The model architecture is similar to CLIP [Rad+21], except that the text encoder is a pre-trained PubMedBERT [Gu+21] model with WordPiece tokenizer [Wu+16]. The model uses ViT-B/16 [Dos+21] as the visual data encoder. During pre-training, the model adopts a contrastive learning approach, and to mitigate memory usage, it utilizes the sharding contrastive loss [Che+22]. For adaptation to VQA, the model incorporates the METER [Dou+22] framework. This involves deploying a Transformer-based co-attention multimodal fusion module that produces cross-modal representations. These representations are then fed into a classifier for the final prediction of answers. The model is evaluated on VQA-RAD [Lau+18] and SLAKE (English) [Liu+21] datasets (see Table 3).
@@
  • UniXGen (Unified chest X-ray and report Generation model) [Lee+23] is a unified model that can generate both reports and view-specific X-rays. The model tokenizes chest X-rays leveraging VQGAN [ERO21], a generative model that amalgamates generative adversarial networks (GANs) with vector quantization (VQ) techniques. VQGAN employs an encoder to transform input images into continuous representations, subsequently using vector quantization to discretize them into learnable codebook vectors. Additionally, VQGAN incorporates a decoder, translating these discrete codes back into images during the generation process. For chest X-rays, multiple views from the same study are tokenized into sequences of discrete visual tokens, demarcated by special tokens to distinguish perspectives. In the case of radiology reports, the model uses a byte-level BPE [WCG20] tokenizer, augmented with sinusoid positional embedding for enhanced representation. The model is based on the Transformer architecture [Vas+17] with multimodal causal attention mask, ensuring that each position in the sequence attends to all previous positions and not future ones. During training, multiple views of chest X-rays and a report embedding are concatenated randomly and fed into the Transformer. The model is optimized using the negative log-likelihood loss function. The model is trained on $208,534$ studies sampled from the MIMIC-CXR [Joh+19a] dataset. Each study contains at most three chest X-rays representing PA (from back to front), AP (from front to back), and lateral views. UniXGen achieves a BLEU-4 score of $0.050$, and using a CheXpert labeler [Irv+19] a precision score of $0.431$, a recall value of $0.410$, and an F1 score of $0.420$ on MIMIC-CXR dataset.
@@
  • RAMM (Retrieval-Augmented bioMedical Multi-modal pretrain-and-finetune paradigm)
@@
    [Yua+23] is a retrieval-augmented VLM tailored for biomedical VQA. The model uses Swin Transformer [Liu+21a] as the image encoder and PubMedBERT [Gu+21] as the text encoder. The visual and textual features are then fused by the multimodal encoder, a 6-layer Transformer [Vas+17]. The model is pre-trained on the MIMIC-CXR [Joh+19a] and ROCO [Pel+18] datasets along with a newly curated dataset PMCPM (PMC-Patients-Multi-modal), consisting of $398,000$ image-text pairs sampled from PMC-OA [Lin+23] dataset. The model also employs contrastive learning, ITM and MLM and the pre-training objective function of the model is the sum of these three tasks. Using contrastive learning, the model aligns images and texts using the cosine similarity metric. The VQA task is viewed as a classification problem and the model is optimized using the cross-entropy loss function. During model fine-tuning, a novel retrieval-attention module fuses the representations of the image-question input with 4 representations of the retrieved image-text pairs from the pre-trained datasets. This allows the model to focus on relevant parts of the retrieved information when generating answers. The model is evaluated on VQA-Med 2019 [Aba+19], VQA-Med 2021 [Ion+21], VQA-RAD [Lau+18], and SLAKE [Liu+21] datasets (see Table 3).

### v2

@@ 4.2.2 Evaluation Metrics for VQA
Human Evaluation
Human evaluation is valuable for assessing a model’s performance and applies not only to tasks such as VQA but also to RG. Human evaluation can be performed for VQA in various ways. For instance, in [moor2023medflamingo], the human evaluation process of MedFlamingo model employs an application featuring a user-friendly interface. Within this interface, medical experts are empowered to evaluate each VQA problem individually, assigning scores ranging from 0 to 10. The final scores of the few-shot performance are $5.61$ on VQA-RAD, $1.81$ on PathVQA, and $4.33$ on the specifically curated Visual USMLE dataset. In contrast, the scores for zero-shot performance are lower, with $3.82$ on RAD-VQA, $1.72$ on PathVQA, and $4.18$ on Visual USMLE.
Table 2: A list of medical VLMs developed for VQA and RG.
@@
[moon2022medvill] Open-I, VQA-RAD
PubMedCLIP dual No ViT-B/32 or RN50 or + ROCO, SLAKE, GH
[eslami2023pubmedclip] RN50$\times$4 + Transformer VQA-RAD
+ BAN
RepsNet dual No ResNeXt-101 + BERT + + VQA-RAD, Site
[tanwani2022repsnet] + BAN + GPT-2 IU-Xray
@@
UniXGen single No VQGAN + Transformer + MIMIC-CXR GH
[lee2023unixgen]
@@
[yuan2023ramm] + multimodal encoder w/ SLAKE, VQA-RAD,
retrieval-atten. module VQA-Med 2019,
VQA-Med 2021
dual No ALBEF + GH
X-REM (ViT-B/16 + BERT MIMIC-CXR,
[jeong2023xrem] + multimodal encoder) MedNLI, RadNLI
@@
[shu2023vismedalpaca] +LLaMa-7B MedQA, PubMedQA
@@
CXR-RePaiR-Gen + prompt manager CXR-PRO,
[ranjit2023cxrrepairgen] + text-davinci-003 MS-CXR
@@
LLaVa-Med single No ViT-L/14 + projection + PMC-15 + GPT-4, GH
[li2023llavamed] layer + LLaMa-7B VQA-RAD, SLAKE,
PathVQA
single No MedCLIP + linear + + GH
XrayGPT transformation layer MIMIC-CXR
[thawkar2023xraygpt] + Vicuna-7B Open-I
@@
dual Yes + ROCO, MedICaT, GH
MUMC ViT-B/12 + BERT ImageCLEF Caption,
[li2023mumc] + multimodal encoder VQA-RAD, SLAKE
+ answer decoder PathVQA
Med-Flamingo single No ViT-L/14 + perceiver + MTB, PMC-OA, GH
[moor2023medflamingo] resampler + LLaMa-7B VQA-RAD, PathVQA,
Visual USMLE
single No BioViL-T + BERT + + GH
RaDialog + prompt manager MIMIC-CXR,
[pellegrini2023radialog] + Vicuna-7B Instruct
4.3 Medical Models
@@
In this part of the review paper, we provide an overview of existing medical VLMs tailored for VQA and/or RG. The information is organized chronologically based on the first appearance of the model. Our focus is mainly on recently introduced open-source or publicly available models. A summary of these VLMs is presented in Table 2.
4.3.1 Medical Vision Language Learner (MedViLL)
@@
MedViLL can process medical images to generate associated reports [moon2022medvill]. The model employs ResNet-50 [he2016resnet], trained on ImageNet [deng2009imagenet], for extracting visual features $v$. The model also leverages a base BERT [devlin2019bert] embedding layers to extract textual features $t$ from clinical reports, which are initially segmented into a sequence of tokens using a WordPiece [wu2016wordpiece] tokenizer. Both textual and visual features incorporate positional information to capture the spatial relationships and sequential order of elements in the input data. To generate a cross-modal representation, vectors $v$ and $t$, along with special tokens [CLS], [SEP]V, [SEP]L are concatenated in a single vector as follows: $(CLS,v,SEP_{V},t,SEP_{L})$. The cross-modal representations are then fed into the BERT model. The MedViLL is pre-training on two tasks: MLM and ITM. The MLM task employs a bidirectional auto-regressive (BAR) self-attention mask, promoting the integration of image and language features. For MLM, a negative log-likelihood loss function is used. The ITM task encourages learning visual and textual features by predicting matching pairs and employs a loss function based on predictions for matching and non-matching pairs. The model is pre-trained on $89,395$ image-report pairs from the MIMIC-CXR [johnson2019mimiccxr] dataset and then fine-tuned for downstream tasks on $3,547$ pairs from the Open-I [demnerfushman2015openi] dataset, an additional dataset comprising radiographic image-report pairs. Only AP view X-rays are included in the analysis of both datasets. VQA is performed on the VQA-RAD [lau2018vqarad] dataset (see Table 3), including open and close-ended questions, where the output representation of [CLS] is used to predict a one-hot encoded answer. For radiology RG fine-tuning, the model uses a sequence-to-sequence (S2S) mask instead of BAR and generates reports by sequentially recovering MASK tokens. RG is evaluated on MIMIC-CXR [johnson2019mimiccxr] and Open-I [demnerfushman2015openi]. MedViLL achieves a BLEU-4 score of $0.066$, a perplexity value of $4.185$, and using a CheXpert labeler [irvin2019chexpert] an accuracy of $84.1\%$, a precision value of $0.698$, a recall value of $0.559$, and an F1 score of $0.621$ on MIMIC-CXR. Additionally, it achieves a BLEU-4 score of $0.049$, a perplexity value of $5.637$, an accuracy of $73.4\%$, a precision value of $0.512$, a recall value of $0.594$, and an F1 score of $0.550$ on Open-I.
4.3.2 PubMedCLIP
@@
PubMedCLIP [eslami2023pubmedclip] is a CLIP-based [radford2021clip] model pre-trained on ROCO [pelka2018roco] dataset, consisting of over 80K image-caption pairs sourced from PMC articles. The model utilizes a CLIP text encoder, which is based on the Transformer [vaswani2017attention] architecture, and three distinct CLIP visual encoders: ViT-B/32 [dosovitskiy2021vit], ResNet-50, and ResNet-50$\times$4 [he2016resnet]. Following the contrastive learning approach in CLIP, the model generates joint representations by computing cosine similarity between textual and visual features. The pre-training objective involves the computation of cross-entropy loss values for both vision and language. These losses are then averaged to derive an overall loss value. Following pre-training, the model is repurposed as a pre-trained visual encoder for VQA. The visual feature in VQA is the concatenation of the model’s output with a convolutional denoising autoencoder (CDAE) [masci2011cdae] output, an image denoising module. The question is encoded using a GloVe [pennington2014glove] word embedding followed by an LSTM [hochreiter1997lstm]. The image and question features are combined using bilinear attention networks (BAN) [kim2018ban], and the resulting representations are passed through an answer classifier, which is a two-layer feedforward NN. The VQA loss is determined by combining the classification and image reconstruction losses. During the VQA fine-tuning, the SLAKE (English) [liu2021slake] and VQA-RAD [lau2018vqarad] datasets, comprising both open- and close-ended questions, are employed. The model’s effectiveness is evaluated in the context of two existing Medical VQA (MedVQA) methods: Mixture of Enhanced Visual Features (MEVF) [zhan2020mevf] and question-conditioned reasoning (QCR) [liu2023condreason]. The assessment involved replacing the visual encoder component in MEVF and QCR with PubMedCLIP and subsequently evaluating the model’s performance. PubMedCLIP in the QCR framework achieves better accuracies on VQA-RAD and SLAKE datasets than in the MEVF framework. The highest accuracies of PubMedCLIP in the QCR framework on both datasets are shown in Table 3.
4.3.3 RepsNet
@@
RepsNet is designed for VQA tasks. It can generate automated medical reports and interpret medical images. The model employs a modified version of the pre-trained ResNeXt-101 [xie2016resnext] as its image encoder and utilizes pre-trained BERT [devlin2019bert] as the text encoder, with text tokenization done through WordPiece [wu2016wordpiece]. Fusion of image and question features is achieved using BAN [kim2018ban]. To align images with textual descriptions, the model employs bidirectional contrastive learning [chen2020contlearningvisrepr]. For VQA tasks, the model is fine-tuned and evaluated on VQA-RAD [lau2018vqarad] (see Table 3). In contrast, for RG, fine-tuning and evaluation are done using IU-Xray [demnerfushman2015openi] dataset. The model categorizes answers through classification for close-ended questions and generates answers using the modified version of GPT-2 language decoder based on image features and prior context. The BLEU-2 and BLEU-4 scores of RepsNet on the IU-Xray dataset are $0.44$ and $0.27$, respectively.
4.3.4 BiomedCLIP
@@
BiomedCLIP is pre-trained on the specifically curated PMC-15 dataset that consists of $15$ M figure-caption pairs derived from the PMC articles [zhang2023biomedclip]. However, the models is not publicly available. The model architecture is similar to CLIP [radford2021clip], except that the text encoder is a pre-trained PubMedBERT [gu2021pubmedbert] model with WordPiece tokenizer [wu2016wordpiece]. The model uses ViT-B/16 [dosovitskiy2021vit] as the visual data encoder. During pre-training, the model adopts a contrastive learning approach, and to mitigate memory usage, it utilizes the sharding contrastive loss [cherti2022shconrtloss]. For adaptation to VQA, the model incorporates the METER [dou2022meter] framework. This involves deploying a Transformer-based co-attention multimodal fusion module that produces cross-modal representations. These representations are then fed into a classifier for the final prediction of answers. The model is evaluated on VQA-RAD [lau2018vqarad] and SLAKE (English) [liu2021slake] datasets (see Table 3).
4.3.5 Unified chest X-ray and report Generation model (UniXGen)
@@
UniXGen is a unified model that can generate both reports and view-specific X-rays [lee2023unixgen]. The model tokenizes chest X-rays leveraging VQGAN [esser2021vqgan], a generative model that amalgamates generative adversarial networks (GANs) with vector quantization (VQ) techniques. VQGAN employs an encoder to transform input images into continuous representations, subsequently using vector quantization to discretize them into learnable codebook vectors. Additionally, VQGAN incorporates a decoder, translating these discrete codes back into images during the generation process. For chest X-rays, multiple views from the same study are tokenized into sequences of discrete visual tokens, demarcated by special tokens to distinguish perspectives. In the case of radiology reports, the model uses the byte-level BPE [wang2020bbpe] tokenizer, augmented with sinusoid positional embedding for enhanced representation. The model is based on the Transformer architecture [vaswani2017attention] with a multimodal causal attention mask, ensuring that each position in the sequence attends to all previous positions and not future ones. During training, multiple views of chest X-rays and a report embedding are concatenated randomly and fed into the Transformer. The model is optimized using the negative log-likelihood loss function. The model is trained on $208,534$ studies sampled from the MIMIC-CXR [johnson2019mimiccxr] dataset. Each study contains at most three chest X-rays representing PA (from back to front), AP (from front to back), and lateral views. UniXGen achieves a BLEU-4 score of $0.050$, and using a CheXpert labeler [irvin2019chexpert] a precision score of $0.431$, a recall value of $0.410$, and an F1 score of $0.420$ on MIMIC-CXR dataset.
@@
4.3.6 Retrieval-Augmented bioMedical Multi-modal Pretrain-and-Finetune Paradigm (RAMM)
@@
RAMM is a retrieval-augmented VLM tailored for biomedical VQA [yuan2023ramm]. The model uses Swin Transformer [liu2021swintransformer] as the image encoder and PubMedBERT [gu2021pubmedbert] as the text encoder. The visual and textual features are then fused by the multimodal encoder, a 6-layer Transformer [vaswani2017attention]. The model is pre-trained on the MIMIC-CXR [johnson2019mimiccxr] and ROCO [pelka2018roco] datasets along with a newly curated PMC-Patients-Multi-modal (PMCPM) dataset, consisting of $398,000$ image-text pairs sampled from PMC-OA [lin2023pmcclip] dataset. The pre-training objective function of the model is the sum of three tasks: contrastive learning, ITM, and MLM. Using contrastive learning, the model aligns images and texts using the cosine similarity metric. The VQA task is viewed as a classification problem, and the model is optimized using the cross-entropy loss function. During model fine-tuning, the retrieval-attention module fuses the representations of the image-question input with four representations of the retrieved image-text pairs from the pre-trained datasets. This allows the model to focus on relevant parts of the retrieved information when generating answers. The model is evaluated on VQA-Med 2019 [abacha2019vqamed19], VQA-Med 2021 [ionescu2021vqamed21], VQA-RAD [lau2018vqarad], and SLAKE [liu2021slake] datasets (see Table 3).
4.3.7 Contrastive X-Ray REport Match (X-REM)